OpenStudy (anonymous):

Simplify the product using the distributive property. (3h - 7)(3h - 6)

OpenStudy (anonymous):

@EchaSetiaWijaya

OpenStudy (unklerhaukus):

\[(3h - 7)(3h - 6)\] \[=(3h\times 3h)+(3h\times-6)+(-7\times3h)+(-7\times-6)\]

OpenStudy (anonymous):

A. 9h2 + 3h - 42 B. 9h2 - 3h - 42 C. 9h2 - 39h + 42 D. 9h2 + 39h + 42

OpenStudy (unklerhaukus):

one of those is correct

OpenStudy (anonymous):

B??

OpenStudy (anonymous):

\[(3h-7)(3h-6)\]\[3h(3h-6)-7(3h-6)\]\[9h ^{2}-18h-21h+42\]\[9h ^{2}-39h+42\]

OpenStudy (anonymous):

the answer is C